Grade 10Differential Calculus

the equation of tangent drawn to the curve y^2-2x^3-4y+8=0 from the point (1,2)is

Matt Farrel
12 Years agoGrade 10

1 Answer

Arun Kumar
ApprovedApproved Tutor Answer12 Years ago
Hello Student,

2ydy/dx-6x^2-4dy/dx=0
=>dy/dx(2y-4)=6x^2
=>dy/dx=3x^2/(y-2)
m=(dy/dx) at (1,2)=infinite
=> x=k
it passes through (1.2)
so 1=k
=> x=1

Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty